OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. Answer y-1=6(x+5) y+5=6(x-1) y+1=6(x-5) y-5=6(x+1)

OpenStudy (anonymous):

Use:\[y - y _{1} = m(x - x _{1})\]since that is the point-slope formula for a line with "m" slope going through (x1, y1)

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

Careful with what is x and what is y.

OpenStudy (anonymous):

`a

OpenStudy (anonymous):

The formula is specific about "+" and "-" signs. And when you subtract a negative, it's like adding a positive.

OpenStudy (anonymous):

I suggest you write out your work so I know where you are having a problem.

OpenStudy (anonymous):

its C.

OpenStudy (anonymous):

because the formula is y-y1=m(x-x1)